OpenStudy (anonymous):

sinx + sin2x + sin3x = cosx + cos2x + cos 3x. find x

OpenStudy (anonymous):

do u know what CosA+cosB =? and what SinA+sinB =?

OpenStudy (anonymous):

i know, but my teacher said that i had to use problem of complex numbers,

OpenStudy (anonymous):

iam getting 5 values for x

OpenStudy (anonymous):

one of them is coming out to be \[\huge \frac{ 2 }{ 3 }(3\pi n-\pi)\]

OpenStudy (anonymous):

what i got is \[\cos(x) = -\frac{ 1 }{ 2 }\]and \[\sin{2x} = \cos{2x}\] Just two equatiions @deena123 what did u get?

OpenStudy (anonymous):

iam getting 5 solutions for x !

OpenStudy (anonymous):

I just got two as you can see ?. what are other three? Can u write your procedure please?

OpenStudy (anonymous):

is the general formula for this question--->\[\huge \frac{ \pi }{ 8 }+n \frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

2 sin2x .cosx +sin 2x = 2 cos2x. cosx + cos2x sin2x(2 cosx+ 1) = cos2x( 2 cosx + 1) sin2x = cos 2x sin2x/cos2x =1 tan2x = 1 2x = tan^-1 (1) 2x =pi/4 x = pi/8. ( a particular solution) hence general solution : x = pi/8 +n pi/2.